# FERGI: Automatic Scoring of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### v3

@@ 6.2.1 Pre-trained Scoring Models
Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to estimate human preferences of text-to-image generation based on the input text prompts and the generated images.
@@ 6.2.2 FAU-Net Valence Score
We aim to use the estimated AU activation values to automatically score user preferences between images generated from the same input text prompt and propose that this scoring more directly reflects the user’s reaction to the generated image and is complementary to the pre-trained scoring models. To achieve this goal, we train an FAU-Net (Facial Action Units Neural Network) to do automatic scoring. In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single ouput neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values. The sigmoid function is used as the activation function for all neurons. The network is trained with a RankNet loss function using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$. For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 minutes.
@@ 6.2.3 Integrate FAU-Net Valence Score with Pre-trained Scoring Models
We separately investigated the effect of integrating the FAU-Net valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Eq. 3 are not modified at this stage.
@@
(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
Figure 5: Weights of hidden nodes. Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights). The hidden nodes are ordered by the weights from them to the output node in this figure.
@@
The results are shown in Tab. 2. Integration with FAU-Net valence score improves the performance of all four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Fig. 4. The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model). We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping). In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
@@ 6.2.5 FAU-Net Analysis
We analyze the weights of the trained FAU-Net to better understand the model. Figure 5 shows the weights from the input preprocessed AU activation values to the nodes in the hidden layer (given that the preprocessings of activation values are still all increasing functions after model training) with the subcaptions illustrating the weight from the corresponding hidden node to the output node. Generally, for hidden nodes positively affecting the output score, they receive positive weights from the activations of AU2, AU6, AU12, AU17, AU20, and AU26 and negative weights from the activations of AU4, AU5, AU9, AU17, and AU25, and for hidden nodes negatively affecting the output score, they receive weights of the opposite signs from the AUs. Similar trends are observed for the weights from the input preprocessed AU activation values and the output node as shown in Fig. 6.
@@ 7 Discussion and Conclusion
Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as the estimation of the user preference. Integrating the FAU-Net valence score with pre-trained scoring models improves their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al. [79] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4. (see Tab. 4 and Fig. 6 of [79]).
We also demonstrated the complementarity between our FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy. As they are respectively better at annotating different sets of image pairs, it is also potentially feasible to use them to fine-tune each other on data where one model gives larger absolute difference of scores for the two images.
In conclusion, we have demonstrated the feasibility of automatically scoring user preferences for image generation from facial expression reaction. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [57], image inpainting [22], and super-resolution [74].
@@ 8 Acknowledgments
We thank Xiaojing Xu and Yuan Tang for helpful prior work and Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions. We are grateful for support from NSF IIS 1817226 and seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion as well as hardware funding from NVIDIA, Adobe, and Sony.
@@ A Supplementary Visuals
To better illustrate the idea of our paper, we present Fig. S1 showing an example of a strong AU4 activation in response to a low-quality image generation and Fig. S2 an example of a strong AU12 activation in response to a high-quality image generation.
@@ B Data Collection
This section provides more details of the collection of the FERGI dataset not provided in Sec. 3. Figs. S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
@@ B.1 Participants
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). and received academic credits as compensation. They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
@@ B.2 Prompt Creation
There are two forms of sessions, structured input (SI) sessions and free-form input (FFI) sessions. Most participants completed the same number of structured input sessions and free-form input sessions. The number of sessions completed by each participant are shown in Tab. S1. By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically). However, the time consumption highly varies for each participant because of numerous reasons (e.g. internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
@@ B.2.1 Structured Input
  • Animate Objects (Fig. 3(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc. Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
  • Inanimate Objects (Fig. 3(b)): Any object that cannot do things, including plants, vehicles, furniture, etc. Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
  • Interactional Relations (Fig. 3(c)): Relationship between an animated object and any other object (something an animate object does to another object). An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Positional Relations (Fig. 3(d)): Relationship between any two objects. An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Other Inputs (Fig. S4): The location of image as background or general environment, the style of the image, and a list of keywords to append at the end of the prompt.
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S3 and S4.
@@ B.2.2 Free-Form Input
The purpose of free-form input is to provide freedom and flexibility for participants to experiment with any prompts to complement the less flexible input style of structured input. The participants are provided with an example prompt and a text box to enter the entire prompt as the final prompt. A screenshot of an example of the prompt creation page for a free-form input session is shown in Fig. S5.
@@ B.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
@@ B.3.1 Image Annotation
A screenshot of an example of the image annotation page (including the details of the questions) is shown in Fig. S8.
@@ B.3.2 Image Ranking
After the participants finish annotating all 5 generated images, they will be asked to rank them from best to worst. All 5 images are displayed side by side, with the left side labeled “Best” and right side labeled “Worst”, and the participants provide ranking by dragging the images to reorder them. A screenshot of an example of the image ranking page is shown in Fig. S9.

### v4

@@ 6.2.1 Pre-trained Scoring Models
Researchers have used the CLIP score [60], Aesthetic score [65], BLIP score [43], ImageReward score [80], PickScore [35], and HPS v2 score [78] to estimate human preferences of text-to-image generation based on input text prompts and generated images.
@@ 6.2.2 FAU-Net Valence Score
We aim to use the estimated AU activation values to automatically score user preferences between images generated from the same input text prompt and propose that this scoring more directly reflects the user’s reaction to the generated image and is complementary to the pre-trained scoring models. To achieve this goal, we train an FAU-Net (Facial Action Units Neural Network) to do automatic scoring. In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single output neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values. The sigmoid function is used as the activation function for all neurons. The network is trained with a RankNet loss function [5] using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$. For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 min.
@@ 6.2.3 Integrate FAU-Net Valence Score with Pre-trained Scoring Models
We separately investigated the effect of integrating the FAU-Net valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Equation 3 are not modified at this stage.
@@
The results are shown in Table 2. Integration with FAU-Net valence score improves the performance of all four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Figure 4. The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model). We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping). In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
To better understand the technical details, let’s first take Figure 4(d) as an example. In Figure 4(d), we first rank the image pairs in the dataset by the absolute value of the difference in FAU-Net scores for the two images, with the assumption that larger score difference implies higher confidence in predicting that the user prefers the image with higher score in the pair and thus higher accuracy. The x-axis of the figure represents “proportion selected”, which is the proportion of “most confident” image pairs—with left side meaning only the image pairs with highest confidence taken into account while right side meaning all image pairs taken into account. The red line in the figure represents the accuracies of FAU-Net for different proportions of “most confident” image pairs selected, and indeed as expected, the accuracy is clearly higher on the left side for the most confident image pairs, which confirms that our model is well calibrated. Now note that on the left side (low proportion of data selected) of the figure, the red line is obviously higher than the other lines, representing other pre-trained scoring models. This implies that for the image pairs with largest differences in their FAU-Net scores, the choice made by the FAU-Net is significantly more accurate than the choice made by the other models—in other words, for the image pairs our FAU-Net model is very good at, the other models are not good at them.
At first glance, this might seem like a matter of course and not impressive, but this is actually not true for other models. Let’s then take Figure 4(a) as an example. As shown in its subcaption in Figure 4(a), the “most confident image pairs to compare” here are selected by choosing the image pairs with largest differences in their ImageReward scores. We can see that the ImageReward score (represented by the blue line) is well calibrated, but is also almost overlapping with the orange line and the green line—they are highly correlated. This implies that for the image pairs the ImageReward model is very good at, the PickScore and the HPS V2 models are also similarly very good at, so they are less complementary to each other. We can see similar results in Figures 4(c) and 4(b) as well. Overall in all of Figures 4(a), 4(c), 4(b) and 4(d), the curves for all three pre-trained scoring models are highly correlated while our FAU-Net score is not correlated with them.
This demonstrates the uniqueness of our FAU-Net valence score, being very complementary to the other models. We can think about how it works as follows. For the image pairs with small differences in FAU-Net scores (for instance if a user does not show much facial expression difference), the FAU-Net scores will not affect the final score too much in Equation 5, so the other models dominate the decision. However, for the image pairs with large differences in FAU-Net scores, the FAU-Net scores dominate the final score in Equation 5, so our final choice decision is now more based on the FAU-Net model than the other models, and these are exactly the image pairs where our FAU-Net score has much better accuracy than the other models.
@@ 6.2.4 Complementarity between FAU-Net and Pre-trained Scoring Models
(e) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
(f) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
(g) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
(h) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
Figure 5: Weights of hidden nodes. Subfigures show the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption giving the weight from that hidden node to the output node (red/green for negative/positive weights). Four hidden nodes with lowest/highest negative/positive weights to the output node are shown here. Full version with all 16 hidden nodes is shown in the supplemental material.
@@ 6.2.5 FAU-Net Analysis
We analyze the weights of the trained FAU-Net to better understand the model. Figure 5 shows the weights from the input preprocessed AU activation values to the nodes in the hidden layer (given that the preprocessings of activation values are still all increasing functions after model training) with the subcaptions illustrating the weight from the corresponding hidden node to the output node. Generally, for hidden nodes positively affecting the output score, they receive positive weights from the activations of AU2, AU6, AU12, AU17, AU20, and AU26 and negative weights from the activations of AU4, AU5, AU9, AU17, and AU25, and for hidden nodes negatively affecting the output score, they receive weights of the opposite signs from the AUs. Similar trends are observed for the weights directly from the input preprocessed AU activation values and the output node as shown in Figure 6.
@@ 7 Discussion and Conclusion
Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as an estimation of user preference. Integrating the FAU-Net valence score with pre-trained scoring models improves their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: [80] used ImageReward (trained from manually labeled human preferences), with only an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4. (see Tab. 4 and Fig. 6 of [80]).
We also demonstrated the complementarity between FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy.
Admittedly, our study has some limitations: (1) users may be reluctant to turn on their cameras in real life; (2) the participants of our study are explicitly aware that their reaction videos are being recorded, which might affect the spontaneousness of their facial expressions [87, 55].
In conclusion, we have demonstrated the feasibility of automatically scoring user preferences for image generation from facial expression reaction. The application of this method may also be potentially generalizable to other image generation tasks, such as image-to-image translation [58], image inpainting [22], and super-resolution [75].
@@ 8 Acknowledgments
We thank Xiaojing Xu and Yuan Tang for helpful prior work, Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions, and reviewers for helpful comments. We are grateful for support from NSF IIS 1817226 and CRCNS 2208362, seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion, and gift support from an anonymous donor, as well as hardware funding from NVIDIA, Adobe, and Sony.
@@ A Supplementary Visuals
To better illustrate the idea of our paper, we present Figure S1 showing an example of a strong AU4 activation in response to a low-quality image generation, and Figure S2 an example of a strong AU12 activation in response to a high-quality image generation.
@@ B Data Collection
Figure S3: Procedure of one session in data collection. A copy of Figure 1 with additional references to the screenshots of each stage. Firstly, the participant creates an input text prompt. Secondly, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately; at the same time, 5 images are generated from the text prompt using Stable Diffusion v1.4. Then, for each image, the participant goes through a 5-second baseline phase, a 5-second image presentation phase, and an image annotation phase with no time restriction. (During the baseline phase, the participant’s baseline facial expression is recorded, and during the image presentation phase, the participant’s facial expression reaction to the generated image is recorded.) Finally, after all 5 images have been presented and annotated, the participant ranks the 5 generated images from best to worst.
This section provides more details of the collection of the FERGI dataset not provided in Section 3. Figure S3 displays a copy of Figure 1, showing the procedure of one session in data collection with additional references to the screenshots of each stage. Figures S4, S5, S6, S7, S8, S9 and S10 display screenshots of an example of one session in data collection.
@@ B.1 Participants
39 participants [^1] The user IDs are up to 040 as shown in Table S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). and received academic credits as compensation. They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
@@ B.2 Prompt Creation
There are two forms of sessions, structured input (SI) sessions and free-form input (FFI) sessions. Most participants completed the same number of structured input sessions and free-form input sessions. The number of sessions completed by each participant are shown in Table S1. By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically). However, the time consumption highly varies for each participant because of numerous reasons (e.g. internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
@@ B.2.1 Structured Input
  • Animate Objects (Figure 4(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc. Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
  • Inanimate Objects (Figure 4(b)): Any object that cannot do things, including plants, vehicles, furniture, etc. Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
  • Interactional Relations (Figure 4(c)): Relationship between an animated object and any other object (something an animate object does to another object). An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Positional Relations (Figure 4(d)): Relationship between any two objects. An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Other Inputs (Figure S5): The location of image as background or general environment, the style of the image, and a list of keywords to append at the end of the prompt.
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figures S4 and S5.
@@ B.2.2 Free-Form Input
The purpose of free-form input is to provide freedom and flexibility for participants to experiment with any prompts to complement the less flexible input style of structured input. The participants are provided with an example prompt and a text box to enter the entire prompt as the final prompt. A screenshot of an example of the prompt creation page for a free-form input session is shown in Figure S6.
@@ B.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figures S7 and S8.
@@ B.3.1 Image Annotation
A screenshot of an example of the image annotation page (including the details of the questions) is shown in Figure S9.
@@ B.3.2 Image Ranking
After the participants finish annotating all 5 generated images, they will be asked to rank them from best to worst. All 5 images are displayed side by side, with the left side labeled “Best” and right side labeled “Worst”, and the participants provide ranking by dragging the images to reorder them. A screenshot of an example of the image ranking page is shown in Figure S10.